OpenStudy (anonymous):

x^-4x+2=0 solve the quadratic equation by completing the square

hero (hero):

x^2-4x+2 = 0

OpenStudy (anonymous):

?

hero (hero):

You forgot to put the square after x

OpenStudy (anonymous):

aww icic lol thank ya

hero (hero):

x^2-4x = -2 x^2-4x + 4 = -2 +4 (x - 2)^2 = 2 x - 2 = +/- sqrt{2} x = +/- sqrt{2} + 2

OpenStudy (anonymous):

so then if i have i have to pick between the from (x+__)^2 or (x-__)^2 i would pick the minus

OpenStudy (anonymous):

and apply the (x-2)^2=2?

hero (hero):

(x-2)^2 The 2 is only negative because the middle term of x^2-4x +4 is negative

OpenStudy (anonymous):

aww ok ok thank